### v1

# Semantic-guided Search for Efficient Program Repair with Large Language Models

Abstract.
Fixing software bugs is crucial yet demands significant resources from developers. Automated Program Repair (APR) is a promising solution to address this challenging task. The emergence of Large Language Models (LLMs) has opened a new era of LLM-based APR, substantially progressed the APR field further. However, conventional LLM-based Program Repair methods face significant challenges regarding memory inefficiency, hindering their scalability and effectiveness. This is largely due to the beam search utilized in the patch generation phase of LLM-based Program Repair, which often requires large beam sizes to search for more potentially good repair candidates.
In this paper, we first show that increases in beam size of even just small-sized LLM (1B-7B parameters) require an extensive GPU resource consumption, leading to up to 80% of recurring crashes due to memory overloads in LLM-based APR. Seemingly simple solutions to reduce memory consumption are (1) to quantize LLM models, i.e., converting the weights of a LLM from high-precision values to lower-precision ones. and (2) to make beam search sequential, i.e., forwarding each beam through the model sequentially and then concatenate them back into a single model output. However, we show that these approaches still do not work via both theoretical analysis and experiments.
To address this, we introduce FLAMES, a novel LLM-based APR technique that employs semantic-guided patch generation to enhance repair effectiveness and memory efficiency. Unlike conventional methods that rely on beam search, FLAMES utilizes greedy decoding to enhance memory efficiency while steering the search to more potentially good repair candidates via a semantic-guided best-first search algorithm. At each decoding step, FLAMES uses semantic feedback from test validation such as the number of passing and failing test cases to select the most promising token to explore further. Our empirical evaluation on the Defects4J and HumanEval-Java datasets shows that FLAMES not only substantially reduces memory consumption by up to 83% compared to conventional LLM-based APR, but also accelerates the repair process. Remarkably, FLAMES successfully generated 133 and 103 correct fixes for 333 and 163 bugs in the Defects4J and HumanEval-Java datasets, respectively. This suggests that FLAMES is not only more efficient but also outperforms state-of-the-art techniques, fixing at least 10 and 11 more bugs than SOTA baselines in the Defects4J and HumanEval-Java datasets, respectively.
Program Repair, Efficiency, Large Language Models
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software testing and debuggingccs: Computing methodologies Natural language processing
1. Introduction
Fixing software bugs is a complex and time-consuming task for developers (Winter et al., 2023). Automated Program Repair (APR)(Goues et al., 2019; Gao et al., 2022) has emerged as a promising solution to alleviate this burden by automatically repairing bugs. Over the past two decades, APR has seen significant advancements (Kim et al., 2013; Mechtaev et al., 2016; Le et al., 2016, 2017; Liu et al., 2019; Ghanbari et al., 2019; Chen et al., 2021a; Koyuncu et al., 2020; Le-Cong et al., 2021; Zhu et al., 2021; Xia and Zhang, 2022; Ye et al., 2022b), with practical applications in the software industry (Winter et al., 2022; Kirbas et al., 2021; Marginean et al., 2019; Bader et al., 2019). Recently, the rise of Large Language Models (LLMs), pre-trained on vast datasets, has further advanced APR, opening a new era of LLM-based APR (Silva et al., 2023; Xia et al., 2023b; Jiang et al., 2023a). LLM-based techniques typically adapt models for APR either through fine-tuning on APR-specific datasets (Silva et al., 2023; Jiang et al., 2023a; Huang et al., 2023; Paul et al., 2023; Zhang et al., 2022; Zirak and Hemmati, 2024) or by using prompting with commercial models (Bouzenia et al., 2024; Yin et al., 2024). In this study, we focus on fine-tuning approaches, as prompting often relies on closed-source and commercial models like ChatGPT, which lack transparency, hindering reproducibility in open science and raising data privacy concerns.
Fine-tuning-based approaches typically refine the weights of Code LLMs, such as CodeLLama (Rozière et al., 2023) and Incoder (Fried et al., 2022) by fine tuning these models on APR datasets. During inference, they generate token sequences to construct candidate patches, prioritizing sequences with high probability scores from the fine-tuned model. Traditionally, beam search, a widely-used search algorithm from natural language processing (Sutskever et al., 2014), is used to optimize this process by maintaining only the top-$n$ nodes at each decoding step, where $n$ is the beam size. A larger beam size provides a more precise approximation to exact decoding, i.e., thoroughly exploring all possible sequences, thereby enhancing the quality of the outputs. This hyper-parameter is even more important in the context of APR, which covers a broader range of possible identifiers and a larger search space than natural language (Lutellier et al., 2020; Chen et al., 2021a; Jiang et al., 2021). It has been shown that a substantial beam size is essential for generating adequate candidate patches for optimal results (Jiang et al., 2021; Lutellier et al., 2020; Zhong et al., 2024), and a bigger beam size leads to higher performance of APR techniques that are based on small-sized deep learning models (less than 300M parameters)  (Tufano et al., 2019; Ye et al., 2022a; Zhong et al., 2024).
But, is it true that increasing the beam size is all we need to improve the performance of APR? We revisit the impact of beam size on the effectiveness of LLM-based APR techniques on larger models (1B-7B parameters). Our findings reveal that increasing the beam size boosts the effectiveness of these techniques, but once the beam size is reached a threshold the performance drops significantly due to memory overloads that cause recurring crashes even on state-of-the-art hardware (NVIDIA A100 GPU, equipped with 80 GB of VRAM). For instances, state-of-the art LLM-based APR techniques such as InCoder (Jiang et al., 2023a), CodeGen (Jiang et al., 2023a), and RepairLLama (Silva et al., 2023) produced between 21% and 46% more plausible patches when the beam size was increased from 10 to 25. Ideally, we expect that further increasing the beam size would lead to better performance of the models for APR. However, when beam size is set higher at 50, 100, and 200, our experiments reveal that performance of these models significantly drops.
Our experiments uncover that the above phenomenon is due to the extensive GPU resource consumption, particularly on Video Random Access Memory (VRAM), caused by larger beam sizes. Often, upgrading VRAM to accommodate larger beam sizes requires GPU upgrades, which are expensive, and thus this creates a costly trade-off between memory efficiency and the performance of LLM-based APRs. For instance, achieving a 46% performance improvement with InCoder-1B required 58% more average and 111% more peak VRAM usage on our dataset. Consequently, further performance gains through increasing beam size demand substantial VRAM, often leading to out-of-memory (OOM) crashes even on cutting-edge hardware and thereby unduly reducing the effectiveness of LLM-based APR techniques. Our experiments using the NVIDIA A100 GPU, equipped with 80 GB of VRAM, showed that Incoder-6B crashed on nearly 80% of the evaluated bugs when using a beam size of 200. This resulted in a 60% reduction in performance compared to a beam size of 10. In summary, we conclude that simply increasing the beam size is not a solution as it comes at the cost of chasing for state-of-the-art hardware.
A seemingly straightforward approach to this problem is to reduce the memory usage of LLM-based program repair through engineering efforts. This can be achieved using two common strategies: (1) quantizing LLM models by converting their weights from high-precision to lower-precision values, and (2) making beam search sequential, where each beam is processed individually through the model before being combined into a single output. However, our findings show that while these methods can reduce memory usage, their memory demands still escalate significantly as beam size increases, leading to substantially high out-of-memory (OOM) rates. Consequently, these engineering efforts alone cannot solve the problem. This realization motivates us to develop a patch generation algorithm that can effectively enhance the performance of LLM-based APR techniques without relying on increasing the beam size and compromising memory efficiency.
In this paper, we introduce FLAMES, an efficient LLM-based Program Repair technique using semantic-guided patch generation. Unlike conventional methods that rely solely on a language model’s knowledge and beam search, our approach aim to leverage semantic feedback from test validations to guide LLMs in patch generation. The core idea of FLAMES is to combine LLM-based and search-based APR. This process begins with generating initial patches using a greedy decoding, i.e., beam search with beam size of 1, followed by iterative, semantic-guided searches to refine these solutions. To achieve this, we employ a best-first search algorithm, specifically PG-TD (Zhang et al., 2023a), along with semantic feedback from test validations to guide the patch generation process in LLMs. FLAMES provides three primary advantages in addressing the limitations associated with beam search. First, it significantly reduces VRAM consumption by using greedy decoding, enhancing memory efficiency. Second, the approach scales well as it does not increase VRAM usage while generating more candidate patches. Third, FLAMES facilitates seamless information exchange between the patch generation and validation phases, enabling the efficient and effective exploration of plausible candidate patches.
We conducted an empirical evaluation of our proposed approach, FLAMES, on a dataset including 333 real-world bugs from Defects4J and 163 synthetic bugs from the HumanEval dataset. Our experiments demonstrate that FLAMES successfully repaired 133 and 103 bugs from the Defects4J and HumanEval datasets, respectively. We compared FLAMES against 15 leading APR techniques. Our results indicate that FLAMES substantially outperforms these baselines, achieving at least 10 and 11 more correct fixes in the Defects4J and HumanEval datasets, respectively. We also found that FLAMES’s semantic-guided patch generation substantially outperforms widely-used patch generation strategies in LLM-based APR, including beam search and multiple sampling. The improvement rates range from 43% to 1145% across five different Large Language Models (LLMs) with three different architectures. Additionally, our analysis of memory efficiency showed that our method could reduce VRAM consumption by 42% to 83%, decreasing peak VRAM requirements from over 80 GB to as low as 12.7 GB across various configurations and models. Additionally, FLAMES also significantly boosts the time efficiency of LLMs, often generating plausible patches faster than conventional LLMs with beam search.
In summary, we made the following contributions:
  • We empirically study the impact of beam size on the effectiveness and memory efficiency of five different LLM-based APR techniques, highlighting the challenges of scaling the search space due to extensive memory consumption of the techniques.
  • We introduce a novel approach, namely FLAMES, which fuses LLM-based and search-based APR, utilizing semantics feedback from patch validation to efficiently discover plausible candidate patches.
  • We empirically evaluate the performance of FLAMES against 15 state-of-the-art baselines. FLAMES correctly fixes 133 bugs from the Defects4J dataset and 103 bugs from the HumanEval-Java dataset, surpassing the best baseline by 10 and 11 fixes, respectively. FLAMES also outperforms widely-used patch generation strategies by 53% - 93%. Moreover, FLAMES substantially reduces memory consumption by up to 83% and accelerates the repair process.
2. Motivation Study
Beam size is a critical hyper-parameter in beam search, the fundamental algorithm for patch generation in LLM-based APR. In this section, we evaluate the impact of beam size on the effectiveness and memory efficiency of LLM-based APR techniques using LLMs with 1B to 7B parameters. While prior studies (Jiang et al., 2021; Lutellier et al., 2020; Xia and Zhang, 2022; Wei et al., 2023) showed that a substantial beam size is needed for optimal performance for APR, our experiments reveal that increasing the beam size unduly hinders memory efficiency and subsequently degrades the performance of LLM-based APR even on state-of-the-art hardware. This motivates a new search algorithm, which we propose in Section 3, to improve the efficiency and effectiveness of LLM-based APR. We structure our empirical study with the following two research questions:
  • RQ1: How does beam size affect memory efficiency of LLM-based APR? We measure VRAM usage and the frequency of out-of-memory crashes in LLM-based APR techniques with different beam sizes to understand the impact of this parameter on memory efficiency. We selected beam sizes of 10, 25, 50, 100, and 200, as these are common in prior works (Ye et al., 2022a; Xia et al., 2023b; Xia and Zhang, 2022; Zhu et al., 2021). We did not go beyond beam size of 200 due to hardware constraints.
  • RQ2: How does beam size affect the effectiveness of LLM-based APR? We investigate the impact of beam size on the effectiveness of 5 LLM-based APR techniques by measuring the number of plausible patches. Similarly to RQ1, we also selected beam sizes of 10, 25, 50, 100, and 200.
For both research questions, we also examine the impact of memory reduction techniques, as described in Section 2.1.3, on our findings.
2.1. Experimental Design
2.1.1. Benchmark Dataset.
To address these research questions, we experiment on Defects4J, a well-established benchmark. This dataset was collected by Just et al.(Just et al., 2014) from 17 open-source Java projects, comprising 835 real-world bugs. In this study, following previous works(Ye et al., 2022a; Le-Cong et al., 2024; Xia and Zhang, 2022; Liu et al., 2019; Chen et al., 2021a), we focus on single-hunk bugs, where the patch is limited to a single contiguous chunk of code at a single location. We select this kind of bug as the existing LLM-based APR, which are studied in our work, such as InCoder and CodeGen (Jiang et al., 2023a), are only fine-tuned for single-hunk bugs (refer to next sections for details).In total, we identified 333 single-hunk bugs using the script provided by the original authors of Defects4J.
2.1.2. Studied LLM-based Program Repair techniques
To select the target techniques for our empirical study, we conducted a literature review and identified suitable techniques based on the following criteria: (1) use beam search as the patch generation strategy, (2) employ LLM with at least one billion parameters, (3) provide accessible fine-tuned models for Automated Program Repair.
Based on these criteria, we identified five LLM-based APR techniques from recent works (Jiang et al., 2023a; Silva et al., 2023) for our study including CodeGen-2B (Nijkamp et al., 2022), CodeGen-6B (Nijkamp et al., 2022), InCoder-1B (Fried et al., 2022), InCoder-6B (Fried et al., 2022), and RepairLlama (Silva et al., 2023). The CodeGen and InCoder models are LLMs which has been fine-tuned for Program Repair by Jiang et al.(Jiang et al., 2023a) and shown the best performance among evaluated techniques. Meanwhile, RepairLlama, developed by Silva et al.(Silva et al., 2023), builds on the CodeLLama (Rozière et al., 2023) model, fine-tuned for Program Repair using the QLoRA (Dettmers et al., 2024) method with optimized input and output representations. Note that, as RepairLlama has already used QLoRA and 8-bit quantization, we do not consider their quantization version of this technique in our follow-up experiments. Detailed information of these techniques is provided in our replication package (Authors, 2023).
2.1.3. Memory Reduction techniques.
In this study, we also investigated the potential of memory reduction techniques through engineering efforts and their impact on our findings. These techniques include:
  • Quantization: A widely used technique to reduce the memory usage of LLMs by converting their weights from high-precision (32-bit) to lower-precision values, such as 8-bit or 4-bit. In this work, we apply 4-bit quantization, reducing the LLM’s weights from 32-bit to 4-bit.
  • Sequential Beam Search: An implementation variant of beam search proposed by Saibo Geng (developers, 2024). This approach processes each beam sequentially through the model before concatenating them into a single output.
2.1.4. Implementation Details.
We implement LLM-based APR techniques using Python, leveraging the PyTorch and HuggingFace libraries for efficient execution. To enhance the accuracy of our implementations and minimize potential bias, we collected and integrated the inference code, including repair prompts and trained models, from their original repositories into a unified framework. All implementations utilize HuggingFace’s standard beam search, a widely used method in LLM-based techniques. We perform inference with a batch size of one, i.e., trying to repair one buggy program at a time. All experiments are conducted on a single NVIDIA H100 GPU with 80GB of VRAM, a state-of-the-art and widely-used hardware for Deep Learning and Large Language Models.
2.2. RQ1: Impact of Beam Size on Memory Efficiency
Figure 1. Memory usage of LLM-Based APR techniques across various beam sizes on an NVIDIA A-100 GPU with 80 GB VRAM, using both original and quantized Models. ‘O‘ and ‘Q‘ denote original and quantized models, respectively. ‘BS‘ and ‘SeqBS‘ denote beam search and sequential beam search, respectively. RepairLlama with quantization and sequential beam search were not assessed because (1) the original models were already quantized and (2) sequential beam search does not support this architecture.
Figure 1 illustrates the experimental results of RQ1. Overall, there is a significant increase in VRAM usage as the beam size expands from 0 to 200. For instance, in CodeGen-2B, average memory usage jumps from approximately 20GB at a beam size of 10 to around 70GB at a beam size of 200. It is important to note that our figures are capped at 80GB due to hardware limitations, representing only the lower bound of the actual numbers, which may be much higher. Similarly, peak memory usage climbs dramatically from about 30GB to 80GB, which is the maximum our hardware can support. These trends are consistent across other LLM-based APR techniques. Notably, this increase in VRAM usage induces a substantial rise in crashes due to OOM errors. For example, the OOM ratio in CodeGen-2B increases from nearly negligible at smaller beam sizes to about 60% at a beam size of 200. Larger models, such as CodeGen-6B and InCoder-6B, show even higher OOM crash rates, exceeding 80
Next, we explore the potential of 4-bit quantization and sequential beam search, which are widely used to optimize model memory requirements, to reduce VRAM usage. Figure 1 shows that quantization significantly reduces both average and peak VRAM usage, decreasing the frequency of OOM crashes. However, sequential beam search does not notably reduce the memory usage of LLM-based APR techniques and even increases the frequency of OOM crashes. Sequential beam search involves trade-offs and does not guarantee reduced memory usage in all scenarios. This observation has been also confirmed by the authors of sequential beam search. Further details of this can be found in our online Appendix (Authors, 2023).
More importantly, we also observe that VRAM usage for LLM-based APR techniques with both quantization and sequential beam search still increases substantially as beam size increases. For instance, average VRAM usage for CodeGen-2B rises from less than 10GB at a beam size of 10 to around 60GB at a beam size of 200. Its peak memory usage also quickly approaches the hardware limit of 80GB. Consequently, even with memory reduction techniques, LLM-based APR techniques continue to experience a significant number of OOM crashes in our benchmark dataset, ranging from 30% to 75% at a beam size of 200.
Answers to RQ1: Both average and peak memory usage of LLM-based APR techniques rise significantly as the beam size increases, inducing up to 80% out-of-memory crashes, on a NVIDIA A100 with 80GB VRAM. Even with memory reduction techniques, memory usage still escalates and crash rates remain substantial as the beam size increase.
2.3. RQ2: Impact of Beam Size on Effectiveness
Table 1. Effectiveness of LLM-based APR techniques was evaluated across different beam sizes on an NVIDIA A100 GPU with 80 GB VRAM, using both original and quantized models. ‘O‘ and ‘Q‘ denote original and quantized models, respectively. ‘BS‘ and ‘SeqBS‘ denote beam search and sequential beam search, respectively. RepairLlama with quantization and sequential beam search were not assessed because (1) the original models were already quantized and (2) sequential beam search does not support this architecture.
Model O + BS Q + BS O + SeqBS
10 25 50 100 200 10 25 50 100 200 10 25 50 100 200
CodeGen-2B 66 80 87 75 38 64 80 98 107 89 65 73 80 70 27
CodeGen-6B 72 93 75 39 11 65 95 114 91 55 62 91 74 39 5
InCoder-1B 74 108 119 120 96 71 102 124 134 126 69 104 107 109 65
InCoder-6B 77 95 92 61 27 77 100 123 114 83 75 88 76 34 14
RepairLlama 139 169 160 134 75 _ _ _ _ _ _ _ _ _ _
Table 1 present our experimental results for RQ2. We can see that the number of plausible patches, only shows an upward trend at smaller beam sizes and remarkably declines at larger sizes. For instance, the number of plausible patches generated by RepairLlama increases from 139 at a beam size of 10 to 169 at 25, then drops significantly to 134 at 100 and further to 75 at 200, although the ratio of plausible patches remains at 0.60. This reduction is primarily due to the high frequency of OOM crashes, which range from 30-60% at larger beam sizes, as depicted in Figure 1. These results highlight the significant impact of LLMs’ extensive memory usage on their performance with large beam sizes.
Next, similar to RQ1, we examine the impact of 4-bit quantization and sequential beam search on memory usage and the effectiveness of LLM-based APR techniques. Overall, the effectiveness of LLM-based APR techniques remains consistent with sequential beam search, maintaining the same trends. Meanwhile, quantization sustains an upward trend in effectiveness up to a beam size of 100. For example, the number of plausible patches generated by the quantized CodeGen-2B increased from 64 at a beam size of 10 to 107 at a beam size of 100, surpassing the original model by 42.7% at the same beam size. This improvement is due to a reduction in OOM crashes, from 30% to less than 5%. However, despite these gains, the effectiveness of quantized models still declines at larger beam sizes, such as 200, similar to the trend seen in the original models. For instance, the effectiveness of the quantized CodeGen-2B dropped from 114 plausible patches at a beam size of 50 to 91 at 100, and further to 55 at 200, reflecting declines of 20% and 48%, respectively.
Answers to RQ2: As beam size increases, the effectiveness of LLM-based APR generally rises, peaking at mid-range beam sizes, e.g., 25 or 50, before declining at larger sizes due to higher OOM crash rates, particularly in larger models. Even with memory reduction techniques, the APR’s effectiveness still quickly degrades at beam sizes larger than 100.
3. Approach
Our findings in Section 2 highlight that scaling beam size to a certain threshold may help improve the effectiveness of APR at the cost of extensive memory consumption, above which the effectiveness quickly degrades even on state-of-the-art hardware.
To address these limitations, we propose a novel approach, FLAMES, that combines LLM-based and search-based APR for an efficient, semantic-guided patch generation. At high level, FLAMES initially generates candidate patches using greedy decoding (i.e., beam search with a beam size of 1) and then leverages semantic feedback from test validation and a specialized best-first search, namely PG-TD (Zhang et al., 2023a), to iteratively explore potential candidate patches. FLAMES uses a small beam size to keep memory usage low, while allowing it to explore a greater number of potential candidate via the guidance of semantic feedback with best-first search.
Figure 2 illustrates the overall pipeline of FLAMES. The search algorithm of FLAMES guides the patch generation process in LLM-based APR through four main phases: Selection, Expansion, Simulation, and Backpropagation. In this section, we first re-formulate the patch generation task in LLM-based APR as a search problem. Then, we present details of these four stages used to generate candidate patches for a given buggy program.
Figure 2. Overview of FLAMES
3.1. Problem Formulation
We first formulate the patch generation process in APR as a Markov Decision Process (Sutton and Barto, 2018) which contains state, action, transition function and reward function as inspried by prior works (Ellis et al., 2019; Zhang et al., 2023a).
State. The state is a triplet $s=(pp,spec,bp)$, where $bp$ is the buggy program which needs to be fixed, $pp$ is a partial patch for $bp$ represented by a sequence of tokens and $spec$ is a specification which describes desired behaviors of $bp$. The initial state is $(\langle sos\rangle,spec,bp)$ where $\langle sos\rangle$ denotes the start of sequence token in LLMs.
Action. An action $a$ is a token from the vocabulary of fine-tuned LLM for APR. We also have special tokens, i.e., terminal tokens, that indicates that a program is complete. These terminal tokens are pre-defined for each LLM. We define that a partial patch $pp$ is complete if and only if the last token of $pp$ is a terminal token.
Transition function. The transition $T$, takes the set of partially-contructed patch $pp$ and applies the action $a$ as follows:
(1) $T(pp,a)=pp\oplus[a]$
where $\oplus$ denote a sequence concatenation.
Reward. The reward of a state $s=(pp,spec,bp)$ is defined to be either:
  • 0 if $pp$ is a partial patch;
  • $R(pp,spec)$ if $pp$ a complete patch, where $R$ is a function which measure how likely $pp$ statisfies $spec$.
Intuitively, $R$ serves for same purpose as objective function in search-based APR (Le Goues et al., 2011; Le et al., 2016).
3.2. Overview
To efficiently and effectively search for repairs, FLAMES maintains a tree structure where each node represents a token and the path from the root to each node represents a partially constructed patch. Then, FLAMES starts from an empty patch and search for a complete patch with optimal reward. The process includes four stages: (1) selection phase for selecting the most promising partial patch, i.e., branch of tree, to visit using a policy algorithm (Section 3.3); (2) expansion phase for expanding the search tree by adding potentially tokens (Section 3.4); (3) evaluation phase for estimating the reward of current state using simulation (Section 3.5); and (4) backpropagation phase for updating value of partial patches that their potential (Section 3.6). We elaborate details of the stages below.
3.3. Selection
At this stage, our goal is to select the most promising partial patch to explore. To achieve this, we employ a policy algorithm that strikes a balance between exploiting potentially high-reward states and exploring less-visited ones. Specifically, an action (i.e., adding a token) is chosen based on three criteria: (1) the action’s historical reward in relation to the current state is high, indicating strong potential; (2) language models (LLMs) predict that the action is a suitable next token, as reflected by its prediction probability; and (3) the resulting state after taking the action is relatively under-explored.
To implement this approach, we evaluate three commonly-used policy algorithms in Monte Carlo Tree Search (MCTS): UCB (Kocsis and Szepesvári, 2006), and two variants of P-UCB (Silver et al., 2017)—one with fixed weights (Fixed P-UCB) and the other with variable weights (Variable P-UCB). Our results indicate that Variable P-UCB is the most effective algorithm for the APR task, leading us to adopt it as our primary policy algorithm (See Section 4.6 for further details.)
3.4. Expansion
This stage aims to expand the search tree by adding possible next tokens. As the number of possible tokens in vocabulary is huge, a random sampling, which is used in conventional MCTS (Kocsis and Szepesvári, 2006) may likely return in invalid and low-potential tokens. Therefore, FLAMES leverages fine-tuned LLM to suggest top-$k$ next tokens with the highest prediction score given the current state following PG-TD. Note that, as a state can be re-visited multiple times during the tree search, FLAMES also leverages a caching mechanism, namely tree structure caching, to avoid redudant computation. In particular, whenever LLMs make a prediction, the input sequence, i.e., state and top-$k$ next tokens is cached using a hash map. Then, when the algorithm needs LLM to predict top-$k$ next tokens, such information can be reused if it exists in the cache. In this work, we set $k=10$ and further discuss the impact of $k$ on our method in Sections 4.6.2.
3.5. Evaluation
This stage aims to evaluate the potential of the current state, i.e., how likely the partial patch leads to a optimal (complete) patch. Similar to expansion, it is impossible to evaluate the current state using an random simulation as used in conventional MCTS algorithm. Consequently, FLAMES leverages beam search to automatically complete the current state, i.e., a partial patch $pp$, to form a complete patch $cp$. However, it is noteworthy that leveraging the beam search as in original PG-TD still hinders memory efficiency. Therefore, instead of beam search, FLAMES utilizes a greedy search, that is equivalent to a beam search with a beam of size 1, to find the complete patch for current state while keeping low memory consumption.
Then, given the completed patch, FLAMES estimates the potential of current state by calculating a reward function $R(cp,spec)$. Theoretically, $spec$ and reward function $R$ can takes any kinds of specification and a corresponding pre-defined reward function. In this work, following common practices in literature (Gao et al., 2022), we leverage developer-written test cases as our specification and utilize the passing ratio, which is widely-used in search based APR (Gao et al., 2022), as our reward function. More formally, given a program $cp$ and a set of test cases $spec$, we define our reward function as follows: $R(cp,spec)=\dfrac{f_{pass}}{f_{pass}+f_{fail}}$ with $f_{pass}$ and $f_{fail}$ is the number of passing and failing test cases observed when running $cp$ on $spec$.
3.6. Backpropagation
This stage aims to update the value of nodes in the tree for guiding next steps of the search process. Intuitively, the stage allows FLAMES to provide feedback from test validation for guiding the patch generation process of LLMs. Particularly, the reward of a state (calculated based on automatically-completed program) is backpropagated to its parents recursively until the root is reached. For all state and action pair $(s^{\prime\prime},a^{\prime\prime})$ along the path from current state to root node, FLAMES updates their Q-value as follows:
(2) $Q(s^{\prime\prime})=max(Q(s^{\prime\prime}),R(cp,spec))$
where $R(cp,spec)$ is reward value of current state - action pair $(s,a)$ with $cp$ is automatically-completed patch of $s$.
3.7. Implementation Details.
We implemented FLAMES by extending the PG-TD (Zhang et al., 2023a) framework, based on the DynaGym toolkit. It is developed in Python using PyTorch, HuggingFace, and OpenAI Gym. We use RepairLLama, the best-performing model among LLM-based repair techniques, as our default LLM-based repair model. For repair, FLAMES generates up to 200 patches and sets a 30-minute timeout for each bug, as recommended by developers (Noller et al., 2022). Experiments are conducted on an NVIDIA A100 GPU with 80GB VRAM, 250 GB RAM, and a 32-core Intel Xeon CPU at 2.90GHz, running Red Hat Enterprise Linux 9.3 and Java 1.8.0_241. This setup is comparable to those used in related works (Silva et al., 2023; Wei et al., 2023).
4. Empirical Evaluation
4.1. Research Questions
Continuing from our motivation study in Section 2, our empirical evaluation in this section aims to answer the following four research questions:
  • RQ3: How effective is FLAMES? We assess the effectiveness of FLAMES on the Defects4J (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023a) datasets, comparing it to 15 leading APR techniques described in Section 4.2.
  • RQ4: How efficient is FLAMES? We evaluate the time and memory efficiency of FLAMES and compare it to the conventional beam search algorithm.
  • RQ5: How effective is FLAMES’s semantic-guided patch generation? We investigate the effectiveness of FLAMES’s semantic-guided patch generation by implementing our approach with five different LLM-based APR techniques and comparing their effectiveness to that of the original models using two widely-used patch generation strategies: beam search (Ye et al., 2022a; Silva et al., 2023) and multiple sampling (Wei et al., 2023; Xia et al., 2023a).
  • RQ6: How do different configurations affect the performance of FLAMES? Finally, we examine the impact of various key factors on FLAMES’s effectiveness, including reward function, expansion size, and policy algorithm.
4.2. Experimental Setup
4.2.1. Benchmark Dataset.
Similar to our initial motivation study (Section 2.1), we evaluate our method using 333 single-hunk bugs from the Defects4J dataset (Just et al., 2014). We also employ the HumanEval-Java dataset, which includes 163 artificial bugs introduced by Jiang et al.(Jiang et al., 2023a). This dataset, recently developed by injecting bugs into the HumanEval benchmark(Chen et al., 2021c), presents a reduced risk of data leakage, where evaluation data might overlap with training data, compared to Defects4J.
4.2.2. APR Baselines.
To evaluate effectiveness of our approach (RQ3), we compare FLAMES against leading baselines across various APR categories: Template-based, NMT-based, Cloze-based, and LLM-based APR techniques. For LLM-based APR, we evaluate five top models—CodeGen-2B, CodeGen-6B, InCoder-1B, InCoder-6B, and RepairLlama—using their best variants identified in RQ2 with beam sizes from 10 to 200. For Cloze-based APR, we include FitRepair (Xia et al., 2023a), AlphaRepair (Xia and Zhang, 2022), and Repilot (Wei et al., 2023). We also compare FLAMES with NMT-based methods: ITER (Ye and Monperrus, 2024), Recoder (Zhu et al., 2021), KNOD (Jiang et al., 2023b), and RewardRepair (Ye et al., 2022a), as well as template-based approaches: GAMMA (Zhang et al., 2023b), TENURE (Meng et al., 2023), and TBar (Liu et al., 2019). Following standard practices (Jiang et al., 2023a, 2021; Chen et al., 2021a; Wei et al., 2023; Xia and Zhang, 2022), we perform evaluations under the assumption of perfect fault localization, which ensures that variations in fault localization techniques do not affect the results. For LLM-based techniques, we use the same configurations as in Section 3.7. For other techniques, we collect bug fix results from their original papers and replication packages, following common practices in prior works (Jiang et al., 2021; Wei et al., 2023; Xia and Zhang, 2022; Ye et al., 2022a).
4.2.3. Patch Generation Algorithm Baselines.
To evaluate the effectiveness of FLAMES’s semantic-guided patch generation (RQ5), we also compared our approach two well-known patch generation strategies: beam search and multiple sampling. For beam search, we leverage the standard implementations from Huggingface library,widely adopted toolset for training and deploying LLMs. For multiple sampling, we first set temperature of LLM to 1 for maximizing their diversity. Then, we repeat the default patch generation of these models (using beam search) until reaching 200 patches.
4.2.4. Evaluation Metrics and Patch Correctness Assessment.
We assess the effectiveness of APR techniques using a standard metric: the number of correct patches. To identify patch correctness, following prior works (Ye et al., 2022a; Le et al., 2019; Xia and Zhang, 2022; Lutellier et al., 2020; Le-Cong et al., 2024), we conduct a manual evaluation to determine if they are either syntactically or semantically equivalent to developer-written patches. Note that, while manual patch correctness assessment is precise, it is resource-intensive and necessitates considerable human effort. Therefore, due to limited human resources, we restrict the use of the number of correct patches as the sole metric for evaluating RQ3, which compares the effectiveness of FLAMES with state-of-the-art baselines. For other research questions, we use the number of plausible patches as a proxy for effectiveness, following prior research (Jiang et al., 2023a; Silva et al., 2023).
4.3. RQ3: Comparison with Existing Techniques
Table 2 reports the number of correct fixes of our proposed technique, FLAMES, and baseline approaches in the Defects4J and HumanEval-Java dataset.
Table 2. Number of correctly fixed bugs by each tool across Defects4J and HumanEval-Java benchmarks.
Techniques Direction Defects4J HumanEval-Java
TBar Template 46 -
GAMMA Template 87 -
TENURE Template 84 -
RewardRepair NMT 79 -
Recoder NMT 43 -
KNOD NMT 110 -
ITER NMT 55 -
AlphaRepair Cloze 95 -
FitRepair Cloze 123 -
Repilot Cloze 116 -
InCoder-1B LLM 57 82
InCoder-6B LLM 50 71
CodeGen-2B LLM 41 60
CodeGen-6B LLM 49 57
RepairLLama LLM 112 92
FLAMES (Ours) LLM 133 103
4.3.1. Repaired Bugs.
Overall, our approach, FLAMES, successfully generates correct fixes for 133 out of 333 single-hunk bugs in Defects4J, outperforming all existing APR techniques. Notably, it corrects 21 more bugs than RepairLlama, the best-performing LLM-based method, which utilizes the same core LLM as FLAMES. This superior performance is primarily due to the memory limitations of RepairLlama, which frequently crashes due to OOM errors with larger beam sizes, as discussed in Section 2.2. Consequently, it cannot scale to higher beam sizes, with the best-performing model capped at a beam size of 25. This limitation results in a lower number of validated patches of 25. In contrast, FLAMES can generate up to 200 patches, addressing scalability issues and boosting repair success rates.
Moreover, our proposed technique also outperforms the leading cloze-based approach, FitRepair, and the top NMT-based technique, KNOD, by repairing 10 and 23 more bugs respectively, corresponding to improvement rates of 8% and 21%. It is noteworthy that FLAMES achieves this performance with a budget of only 200 validated patches and a 30-minute timeout for each bug, while FitRepair and KNOD require 1000-5000 validated patches and a 5-hour timeout for each bug.
4.3.2. Unique Repaired Bugs.
Next, to further demonstrate the effectiveness of our approach, we evaluate the number of unique bugs that only FLAMES can fix. Figure 3 illustrates our uniqueness analysis on Defects4J dataset of our approach with respects to the best baselines for NMT-based (KNOD (Jiang et al., 2023b)), Cloze-based (FitRepair (Xia et al., 2023a)) and LLM-based (RepairLLama (Silva et al., 2023)) APR approaches.
Figure 3. Unique Repaired Bugs in Defects4J
From Figure 3, we can see that FLAMES can fix 14 unique bugs in Defects4J dataset, which were not fixed by any state-of-the-art baselines. Listing 1 illustrates the unique correct patches generated by FLAMES for the Cli-4 in the Defects4J dataset. These patches involve highly complex additional code, including intricate if-else statements, making them challenging to generate with an insufficient patch space. Consequently, LLM-based approaches like RepairLLama, which are limited by memory constraints, fail to produce correct patches for these bugs.
Furthermore, although cloze-based, NMT-based, and template-based approaches can generate a large number of patches (ranging from 1,000 to 5,000), they still fail to produce the correct ones. We believe this is because these methods rely solely on the models’ output probabilities during patch generation, resulting in inefficiencies when searching for plausible patches. For example, FitRepair generates 5,000 patches and validates the top 1,000 within 5 hours, yet it still fails to produce correct patches for the Cli-4 bug. In contrast, FLAMES incorporates feedback from validation directly into the patch generation process using PG-TD, offering valuable guidance for creating correct patches. As a result, FLAMES can generate patches for these complex bugs with only 200 validated patches and a 30-minute timeout.
Additionally, we also found that FLAMES achieve the highest number of unique fixes among baselines. Particularly, RepairLlama, KNOD, and FitRepair contribute 6, 5, and 9 unique fixes, respectively while FLAMES can fix 14 unique bugs. Moreover, there are 85 correct fixes that are not addressed by these four leading techniques. By integrating all techniques, we successfully fix 299 bugs. This demonstrates that FLAMES can be combined with other methods to significantly increase the number of correct fixes generated.
4.3.3. Generalizability.
Finally, to address concerns of data leakage (Jiang et al., 2023a) and benchmark overfitting (Durieux et al., 2019), we extend our evaluation of FLAMES to include HumanEval-Java. Due to the lack of data on existing techniques in the literature, our analysis primarily compares the effectiveness of FLAMES with LLM-based approaches. Detailed results are presented in Table 2.
Listing 1: FLAMES’s correct patch for the bug Cli-4 in Defects4J dataset
if (requiredOptions.size() > 0)
Iterator iter = requiredOptions.iterator();
+ StringBuffer buff = new StringBuffer("Missing required option");
+ if (requiredOptions.size() > 1)
+ buff.append("s");
+ buff.append(": ");
- StringBuffer buff = new StringBuffer();
FLAMES achieves the highest success rate, correctly fixing 103 out of 164 bugs in the HumanEval dataset. Specifically, FLAMES fixes 11 more bugs than the best-performing baseline, RepairLlama, which generates 92 correct fixes, showing an improvement of 12%. Compared to other models, FLAMES is 26% and 45% more effective than InCoder-1B and InCoder-6B, respectively. This performance gap widens when compared to the CodeGen variants; FLAMES outperforms CodeGen-2B and CodeGen-1B by approximately 72% and 81%, respectively.
Answers to RQ3 (Effectiveness): FLAMES successfully fixes 133 and 103 bugs in the Defects4J and HumanEval-Java benchmarks, outperforming current state-of-the-art APR techniques. Additionally, FLAMES uniquely fixes 14 bugs in the Defects4J dataset that cannot be fixed by existing APR techniques.
4.4. RQ4: Efficiency
4.4.1. Memory efficiency.
Table 4 compares the memory usage of FLAMES with three baselines: (1) the original LLM with beam search, (2) the quantized LLM with beam search, and (3) the original LLM with sequential beam search, across five different LLMs. We evaluate VRAM utilization using three key metrics: average VRAM usage, peak VRAM usage observed in our dataset, and out-of-memory (OOM) ratios. Since our hardware is capped at 80GB VRAM, we assign a memory usage of 80GB to data points that experience OOM errors. Consequently, in cases of OOM, the reported memory usage represent a lower bound of the actual values.
Table 3. Comparison of memory usage among the original LLM (Original), quantized LLM (Quantized), original LLM with sequential beam search (SeqBS) and FLAMES.
Models Method Average Peak OOM
(GB) (GB) (%)
Original 72.2 80 62.8
CodeGen-2B Quantized 52.9 80 28.5
SeqBS 68.2 80 69.1
Flames 12.1 13.9 0.0
Original 78.7 80 80.5
CodeGen-6B Quantized 64.0 80 49.8
SeqBS 75.5 80 85.3
Flames 27.1 31.7 0.0
Original 52.7 80 31.5
Incoder-1B Quantized 31.2 80 6.9
SeqBS 56.1 80 49.5
Flames 7.4 12.7 0.0
Original 76.8 80 78.1
Incoder-6B Quantized 57.2 80 39.0
SeqBS 75.6 80 86.6
Flames 32.4 37.4 0.0
Original 66.6 80 62.8
RepairLlama Quantized - - -
SeqBS - - -
Flames 8.1 22.8 0.0
Table 4. Comparison of average running times of FLAMES and beam search (BS) for finding plausible patches with statistical measures on different models and beam sizes
Model Beam Avg. Time (s) $|\mathbf{\delta}|$ p-value
Flames BS
CodeGen-2B 10 37.99 49.43 0.33 0.002
25 55.89 63.12 0.22 0.013
50 68.66 88.07 0.24 0.004
100 92.63 124.72 0.27 0.005
200 93.88 133.40 0.37 0.006
InCoder-1B 10 48.18 68.71 0.35 0.001
25 91.39 86.82 0.07 0.195
50 115.79 138.61 0.11 0.081
100 129.62 177.19 0.12 0.059
200 161.58 199.17 0.08 0.178
RepairLLama 10 44.40 57.55 0.25 0.001
25 90.95 79.21 0.13 0.027
50 103.79 96.81 0.15 0.014
100 126.74 111.42 0.15 0.022
200 98.42 101.35 0.22 0.018
We can see that the average VRAM usage of FLAMES shows a substantial decrease across all models evaluated. For instance, with the CodeGen-2B model, average usage drops from at least 72.2 GB in the original configuration to just 12.1 GB using FLAMES with a reduction of over 83%. This trend of reduced VRAM usage is consistent across other models, such as the Incoder-1B, where average usage is reduced from at least 52.7 GB to 7.4 GB.
In terms of peak VRAM usage, while the baselines frequently reach the 80 GB VRAM limit, FLAMES maintains peak values significantly well below this threshold, such as 13.9 GB for CodeGen-2B and 37.4 GB for Incoder-6B. This effective management of peak VRAM not only avoids bottlenecks and crashes from OOM errors but also enhances the practical usability of these models on limited-resource hardware environments.
As a result, the FLAMES method achieves a zero out-of-memory (OOM) ratio across all models, in contrast to the higher OOM ratios observed with the original configurations. For example, the CodeGen-6B model under the original methodology reports an OOM ratio of 80.5%, which significantly limits model reliability. The absence of OOM errors with FLAMES suggests a notable enhancement in the stability and memory efficiency of LLMs.
Answers to RQ4.1 (Memory Efficiency): FLAMES can significantly reduce VRAM usage in LLMs, successfully minimizing average VRAM usage by up to 83% and reducing peak VRAM requirements from 80 GB to as low as 12.7 GB across various configurations and models. Additionally, FLAMES do not introduce any OOM errors while baselines witness OOMs in up to 86.6% of data points.
4.4.2. Time efficiency.
In this experiment, we evaluate the time efficiency of LLM-based APR using our proposed method, FLAMES, compared to traditional beam search techniques. Our aim is to validate the hypothesis that ”FLAMES can identify plausible patches more quickly than the traditional beam search algorithm”. To validate this hypothesis, we measure the time each method takes to successfully identify plausible patches, employing the Mann-Whitney-Wilcoxon (MWW) statistical test and Cliff’s Delta for effect size assessment. This experiment focuses on the CodeGen-2B, InCoder-1B, and RepairLlama models, chosen for their lower incidence of out-of-memory (OOM) crashes, ensuring a robust sample size for reliable testing. The detailed results are presented in Table 4.
For the CodeGen-2B model, our hypothesis is substantiated across all beam sizes, with p-values ranging from 0.002 to 0.013 and effect sizes between 0.22 and 0.37, indicating that FLAMES is consistently more time-efficient than beam search. Specifically, FLAMES shows lower running times across all beam sizes, reducing time by approximately 11.44 seconds at a beam size of 10 and up to nearly 39.52 seconds at a beam size of 200.
The RepairLlama model also supports the hypothesis with p-values between 0.001 and 0.027 across all beam sizes, confirming the time efficiency of FLAMES over beam search, although the effect sizes range from negligible to small. Consequently, FLAMES demonstrates lower average running times at the beam sizes of 10 and 200. For other beam sizes, while FLAMES generally performs faster than beam search, occasional longer run times by FLAMES skew the average, leading to variability in results.
In contrast, the InCoder-1B model exhibits greater variability between the two methods. While FLAMES tends to generate patches faster at higher beam sizes (50, 100, and 200), it underperforms at a beam size of 25 and shows no significant difference at 200. The most favorable performance for FLAMES is observed at a beam size of 10, where it significantly outperforms beam search (p-value of 0.001). However, for beam sizes 25 and 200, the differences in running times lack statistical significance, indicating that the benefits of FLAMES do not uniformly apply across all configurations within this model.
Answers to RQ4.2 (Time Efficiency): FLAMES significantly improves time efficiency in LLMs, surpassing traditional beam search methods.
4.5. RQ5: Effectiveness of semantic-guided patch generation
Table 5. The number of plausible patches generated by FLAMES’s patch generation algorithm compared to beam search (BS) and multiple sampling (MS) on different LLMs
LLM Beam Search Multiple Sampling FLAMES
CodeGen-2B 38 71 123
InCoder-1B 96 88 138
RepairLLama 75 137 210
CodeGen-6B 11 71 137
InCoder-6B 27 87 144
In this experiment, we assess the effectiveness of FLAMES’s semantic-guided patch generation versus beam search and multiple sampling. We evaluated these methods with various LLMs: CodeGen-2B, InCoder-1B, RepairLlama, CodeGen-6B, and InCoder-6B, measuring effectiveness by the number of plausible patches. Detailed results are in Table 5.
FLAMES consistently outperforms beam search across all models, with improvement rates ranging from 43% to 1145%. For instance, with RepairLlama, FLAMES generates 210 plausible patches, nearly three times the 75 produced by beam search. This is due to FLAMES ’s effective use of GPU resources which allows for a more thorough exploration of the large patch space. In contrast, beam search requires substantial VRAM, resulting in a high out-of-memory crash rate as discussed in RQ4 (Section 4.4).
We also observes that multiple sampling can significantly improve upon beam search across various models. This approach involves using beam search with a low beam size and then sampling the model multiple times to generate additional candidates. As a result, it can partially mitigate the memory inefficiencies of beam search and scale the patch space. Despite this, FLAMES still outperforms multiple sampling across models by 53% to 93%. We suspect that this is because our patch generation process is guided by feedback from test validation, whereas multiple sampling are purely random based on model’s outputs.
In conclusion, the consistent performance improvements across various models strongly support the effectiveness of FLAMES’s semantic-guided patch generation compared to traditional beam search and multiple sampling.
Answers to RQ5 (Effectiveness of Patch Generation): FLAMES’s semantic-guided patch generation demonstrates superior performance across multiple LLMs in APR tasks, consistently outperforming widely-used patch generation methods: beam search and multiple sampling by at least 43%.
4.6. RQ6: Ablation Study
To systematically assess how each component influence the performance of FLAMES, we investigated three key factors: the semantic-guided reward function (for evaluation phase), the size of the expansion $k$ (for expansion phase), and the choice of policy algorithm (for selection phase).
4.6.1. Reward Function.
In this experiment, we evaluate the effectiveness of our reward function, as detailed in Section 3.5. We hypothesize that this function will guide the generation of more plausible patches by leveraging the ratio of passing test cases. To test this hypothesis, we compare our approach with a variant that depends solely on the models’ probability outputs and disregards feedback from our semantic reward function. Our experimental results showed that incorporating a semantic reward function can lead to a 40% increase, from 150 to 210, of plausible patches . This substantial improvement highlights the value of semantic guidance in refining the search process to concentrate on patches that is more likely to succeed.
4.6.2. Expansion Size ($k$).
Next, we explore how the expansion size $k$ affects FLAMES. To investigate this, we conducted experiments with various values of $k$, including 3, 5, 7, and 10. Our findings show that as $k$ increased from 3 to 10, the number of plausible patches rose from 190 to 210. Correspondingly, the average time required per patch also increased, from 49.19 seconds when $k=3$ to 98.41 seconds when $k=10$. This trend suggests that a larger $k$ allows for a more thorough exploration of the solution space, which enhances the likelihood of identifying high-quality patches. However, this comes at the expense of increased computational time, highlighting a trade-off between depth of exploration and performance efficiency.
4.6.3. Policy Algorithms.
Finally, we investigated the impact of different policy algorithms on the generation of plausible patches. We examined the performance enhancement observed with advanced policy algorithms, including the original UCB (Kocsis and Szepesvári, 2006), and two variants of P-UCB(Silver et al., 2017)—one with fixed weights (Fixed P-UCB) and one with variable weights (Variable P-UCB). Our experimental results show that Variable P-UCB is the most effective algorithm, generating 210 plausible patches. This policy algorithm outperforms UCB and Fixed P-UCB by 17% (179 plausible patches) and 11% (188 plausible patches), respectively. The performance of Variable P-UCB likely stems from its ability to dynamically adjust the balance between exploration and exploitation during the search process, thereby optimizing the discovery of plausible patches.
5. Discussion
5.1. Limitations
First, while our approach, FLAMES, correctly fixes 21 more bugs and 42 unique bugs compared to the original RepairLlama, there are 21 bugs that the original model could repair but FLAMES could not. In a deeper analysis, we found that this issue is primarily due to the relatively small expansion size ($k=10$) used in our experiments. This limitation compromise FLAMES’s exploration, resulting in a less comprehensive search space than that of RepairLlama’s beam search, which considers up to 25 possible tokens of expansions. Consequently, FLAMES misses some bugs that could potentially be fixed with a broader search scope. Although a straightforward solution would be to increase the expansion size $k$, doing so introduces a trade-off with time efficiency, as shown in Section 4.6.2. Therefore, further analysis is essential to optimize the balance between expansion size and efficiency to enhance FLAMES’s performance.
Second, FLAMES is constrained by the design of its reward function, which relies on the number of passing and failing test cases. While our experiements demonstrated the effectiveness of our reward function on guiding LLM-based APR to generate more plausible patches and thereby more correct patches. It is important to note that a substantial increase in plausible patches does not always result in a proportionally large increase in correct patches. Moreover, it also raise the problem of test overfitting as observed in search-based APR (Qi et al., 2015; Le et al., 2018). We believe that refining the reward function with additional criteria such as naturalness (Xia et al., 2023c; Kang and Yoo, 2022), syntax/semantic similarity (Xin and Reiss, 2017; Wen et al., 2018) or history versions (Nguyen et al., 2024; Le and Le, 2021; Tan and Roychoudhury, 2015) could enhance both the effectiveness and precision of FLAMES.
5.2. Threats to validity
5.2.1. Internal Validity.
A main threat to internal validity stems from the issue of data leakage, where evaluation data might overlap with training and fine-tuning data. Following prior studies (Silva et al., 2023; Jiang et al., 2023a), we employed the HumanEval-Java dataset for evaluation, which comprises 163 artificially introduced bugs within the HumanEval benchmark (Chen et al., 2021c), thereby minimizing the risk of data leakage. Another concern is the manual assessment of patch correctness. To address this, authors and an external annotator rigorously analyze each patch to confirm semantic equivalence to developer-written patches. Moreover, to ensure transparency, we also have made all patches generated by NPR techniques and our assessment results available in our replication package.
5.2.2. External Validity.
Our research shares a common threat to external validity with prior works (Jiang et al., 2021; Silva et al., 2023; Xia and Zhang, 2022; Xia et al., 2023d; Wei et al., 2023; Le-Cong et al., 2024) that our results may not generalize across different programming languages and datasets. To mitigate this risk, we evaluated our model, FLAMES, and state-of-the-art baselines using two well-known datasets, Defects4J (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023a), and found our performance robust across these datasets. Additionally, the core contributions of our approach are language-agnostic, suggesting potential applicability to various programming languages. Therefore, we believe this risk to be minimal.
5.2.3. Construct Validity.
Our methodology may encounter threats to construct validity concerning the appropriateness of our evaluation metrics. Due to limited human resources, beside the number of correct patches, we leverages the number of plausible patches as as a proxy effectiveness metric, which might not fully capture the actual effectiveness of APR techniques. To minimize this risk, we only utilized this metric when comparing different variants of the same APR techniques, where we observed consistent trends between correct patches and plausible patches.
6. Related Works
6.1. Large Language Models for Automated Program Repair
Recent advancements in Large Language Models (LLMs), pre-trained on extensive datasets, have significantly advanced various coding tasks (Fan et al., 2023), such as code generation (Nijkamp et al., 2022; Fried et al., 2022; Chen et al., 2021c), understanding (Nam et al., 2024; Ahmed and Devanbu, 2022; Zhou et al., 2024), and analysis (Li et al., 2024; Kazerounian et al., 2021; Le-Cong et al., 2022). In the field of APR, LLMs have also shown great promise when employed through either Fine-tuning techniques (Silva et al., 2023; Jiang et al., 2023a; Huang et al., 2023; Paul et al., 2023; Zhang et al., 2022; Zirak and Hemmati, 2024) or Prompting methods (Ahmed and Devanbu, 2023; Xia et al., 2023d; Xia and Zhang, 2023).
Fine-tuning approaches concentrate on refining the weights of code-specific LLMs like CodeLLama (Rozière et al., 2023) and Incoder (Fried et al., 2022), using APR datasets including MegaDiff (Monperrus et al., 2021). Notably, Jiang et al.(Jiang et al., 2023a) demonstrated significant enhancements in LLMs for APR through full-parameter fine-tuning, showing remarkable improvements over the original models. Silver et al.(Silva et al., 2023) explored the effectiveness of the QLORA (Dettmers et al., 2024) fine-tuning method, assessing various input and output formats in LLM-based APR and introducing RepairLLama, an advanced APR technique using CodeLLama. Jin et al.(Jin et al., 2023) introduced InferFix for fixing security vulnerabilities by fine-tuning Codex model and retrieval augmented generation.
Prompting approaches aim to directly leverage LLMs without finetuning. Particularly, they mainly focus on designing effective prompting which provide feedbacks to guide LLMs to generate correct programs. Notably, Xia et al. (Xia and Zhang, 2023) propose conversation-driven APR, namely ChatRepair, which guide ChatGPT with instant feedbacks such as test name, relevant test code, and error message, to perform APR in a conversational style. Ahmed et al. (Ahmed and Devanbu, 2023) and Yin et al. (Yin et al., 2024) levarages chain-of-thought prompting to improve the effectiveness of LLMs on APR. Zhang et al.(Zhang et al., 2024) and Bouzenia et al.(Bouzenia et al., 2024) introduce AutoCodeRover and RepairAgent,tools that enable LLMs to follow human-like debugging steps using multi-agent systems to fix real-world issues.
Different from these studies that typically employ beam search to generate candidate patches, our work empirically investigates the limitations of beam search, particularly when implemented on standard hardware setups. Motivated by these shortcomings, we propose the use of semantic-guided best-first search to guide the repair process of LLMs.
6.2. Efficiency of Automated Program Repair
In addition to evaluating effectiveness, several studies also focus on the efficiency of APR (Liu et al., 2020; Chen et al., 2021b; Qi et al., 2013; Xiao et al., 2024; Chen et al., 2022). Liu et al. (Liu et al., 2020) conducted a systematic assessment of 16 APR techniques, revealing inefficiencies in state-of-the-art APR tools and advocating for further industry exploration of efficiency impacts. Chen et al. (Chen et al., 2021b) introduced UniAPR, a patch validation framework that enhances efficiency by avoiding unnecessary restarts of the JVM for all existing bytecode and source code-level APR techniques. Xiao et al. (Xiao et al., 2023) developed ExpressAPR, which accelerates APR by integrating five mechanisms: mutant schemata, mutant deduplication, test virtualization, test prioritization, and parallelization. Chen et al. (Chen et al., 2022) proposed utilizing XGBoost to develop on-the-fly prioritization techniques to expedite APR techniques. Kim et al. proposed CASINO (Kim et al., 2023) and Gresino (Kim et al., 2024), novel patch-scheduling algorithms for enhancing time efficiency of APR techniques.
Different from these studies that typically focus on time efficiency, our approach focus on memory efficiency, motivated by the extensive memory requirements of LLMs. Our work is closely related RepairLLama (Silva et al., 2023), which explores efficient fine-tuning of LLMs during the training phase. Different from this work, we investigate memory efficiency during the patch generation phase.
7. Conclusion
In this study, we empirically evaluated the impact of beam size on the memory efficiency and effectiveness of LLM-based APR techniques. Our findings reveal that an increase in beam size results in significantly higher VRAM usage. Consequently, increasing the beam size cause numerous out-of-memory crashes and subsequently unduly degrades the performance of LLM-based APR. To address this challenge, we introduced FLAMES —the first APR approach that integrates LLMs with a semantic-guided best-first algorithm to guide the repair process. Our evaluation on the Defects4J and HumanEval-Java datasets demonstrates that FLAMES achieves state-of-the-art results. Additionally, FLAMES significantly reduces memory consumption by up to 83% and accelerates the repair process compared to conventional LLM-based APR techniques.
8. Data Availability.
We published our GitHub repository at https://github.com/thanhlecongg/FLAMES and replication package including dataset and results and appendix at  (Authors, 2023).
Acknowledgement
This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative. Thanh Le-Cong is partially supported by Google through its Ph.D. Fellowship program. Xuan-Bach D. Le is supported by the Australian Government through the Australian Research Council’s Discovery Early Career Researcher Award (DECRA) funding